OpenStudy (anonymous):

A car travels at an average speed of 56 miles per hour. How many miles does it travel in 5 hours and 45 minutes?

OpenStudy (anonymous):

d = r times t

OpenStudy (anonymous):

s=d/t ,s=speed,d=distance,t=time d= s*t here s=56 , t= 5+(45/60) hours

OpenStudy (anonymous):

Answer would be 322 miles. d = 56 x 5.75.

OpenStudy (anonymous):

I thought it was distance / time. but this would make sence. Thank you so much. :)

OpenStudy (anonymous):

sense*